# Quantum Curriculum Learning

### v1

@@ Abstract
Quantum machine learning (QML) requires significant quantum resources to achieve quantum advantage. Research should prioritize both the efficient design of quantum architectures and the development of learning strategies to optimize resource usage. We propose a framework called quantum curriculum learning (Q-CurL) for quantum data, where the curriculum introduces simpler tasks or data to the learning model before progressing to more challenging ones. We define the curriculum criteria based on the data density ratio between tasks to determine the curriculum order. We also implement a dynamic learning schedule to emphasize the significance of quantum data in optimizing the loss function. Empirical evidence shows that Q-CurL enhances the training convergence and the generalization for unitary learning tasks and improves the robustness of quantum phase recognition tasks. Our framework provides a general learning strategy, bringing QML closer to realizing practical advantages.
@@
Introduction.— In the emerging field of quantum computing (QC), there is promise in using quantum computers at sufficiently large scales to solve certain machine learning (ML) problems much more efficiently than classical methods. This synergy between ML and QC has led to the creation of a new field known as quantum machine learning (QML) [1], even though the practicality of its applications remains unclear. There is a question as to whether speed is the only metric by which QML algorithms should be judged [2]. In fact, classical ML focuses on identifying and reproducing features from data statistics. The initial hope is that QML can detect correlations in the classical data or generate new patterns that would be difficult for classical algorithms to achieve [3, 4, 5, 6, 7]. However, there is no clear evidence that analyzing classical data inherently requires quantum effects. This suggests a fundamental shift in the research community’s perspective: it is preferable to use QML on data that is already quantum in nature [8, 9, 10, 11, 12, 13].
@@
The typical learning process in QML involves extensive exploration within the domain landscape of a loss function. This function measures the discrepancy between the quantum model’s predictions and the actual values, aiming to locate its minimum using classical computers. However, the optimization often encounters pitfalls such as getting trapped in local minima [14, 15] or barren plateau regions [16]. These scenarios require substantial quantum computational resources to navigate the loss landscape successfully. Additionally, improving task accuracies necessitates evaluating numerous model configurations, especially against extensive datasets. Given the limitation of current quantum resources, this poses a considerable challenge. Therefore, in designing QML models, we must focus not only on their architectural aspects but also on efficient learning strategies to minimize the use of quantum resources.
@@
Task-based Q-CurL.— We formulate a framework for task-based Q-CurL. The target of learning is to find a function (or hypothesis, prediction model) $h:{\mathcal{X}}\to{\mathcal{Y}}$ within a hypothesis set ${\mathcal{H}}$ that approximates the true function $f$ mapping ${\bm{x}}\in{\mathcal{X}}$ to ${\bm{y}}\in{\mathcal{Y}}$ such that $h({\bm{x}})\approx f({\bm{x}})$. To evaluate the correctness of the hypothesis $h$ given the data $({\bm{x}},{\bm{y}})$, the loss function $\ell:{\mathcal{Y}}\times{\mathcal{Y}}\to\mathbb{R}$ is used to measure the approximation error $\ell(h({\bm{x}}),{\bm{y}})$ between the prediction $h({\bm{x}})$ and the target ${\bm{y}}$. We aim to find a hypothesis $h\in{\mathcal{H}}$ that minimizes the expected risk over the distribution $P({\mathcal{X}},{\mathcal{Y}})$:
@@
Depending on the problem, we can decide the curriculum weight $c_{M,m}$, where a larger $c_{M,m}$ indicates a greater benefit of solving ${\mathcal{T}}_{m}$ for improving the performance on ${\mathcal{T}}_{M}$. We evaluate the contribution of solving task ${\mathcal{T}}_{i}$ to the main task ${\mathcal{T}}_{M}$ by transforming the expected risk of training ${\mathcal{T}}_{M}$ as follows (see [22] for more details):
@@
The curriculum weight $c_{M,m}$ can be determined using the density ratio $r({\bm{x}},{\bm{y}})=\dfrac{p^{(M)}({\bm{x}},{\bm{y}})}{p^{(m)}({\bm{x}},{\bm{y}})}$ without requiring the density estimation of $p^{(M)}({\bm{x}},{\bm{y}})$ and $p^{(m)}({\bm{x}},{\bm{y}})$. The key idea is to model $r({\bm{x}},{\bm{y}})$ using a linear model $\hat{r}({\bm{x}},{\bm{y}}):={\bm{\alpha}}^{\top}{\bm{\phi}}({\bm{x}},{\bm{y}})=\sum_{i=1}^{N_{M}}\alpha_{i}\phi_{i}({\bm{x}},{\bm{y}}),$ where the vector of basis functions is ${\bm{\phi}}({\bm{x}},{\bm{y}})=(\phi_{1}({\bm{x}},{\bm{y}}),\ldots,\phi_{N_{M}}({\bm{x}},{\bm{y}}))$, and the parameter vector ${\bm{\alpha}}=(\alpha_{1},\ldots,\alpha_{N_{M}})^{\top}$ is learned from data (see [22] for the detailed derivation, which is referenced from Ref. [23]).
@@
Data-based Q-CurL.— Next, we present a form of data-based Q-CurL that dynamically predicts the easiness of each sample at each training epoch, such that easy samples are emphasized with large weights during the early stages of training and conversely. Apart from improving generalization, the benefit of data-based Q-CurL lies in its resistance to noise, which is especially needed in QML. Existing QML models can accurately fit partially corrupted labels to quantum states in the training data but fail on the test data [25]. We show that data-based Q-CurL can enhance the robustness based on the dynamic weighting of the difficulty fitting quantum data to corrupted labels.
@@
To evaluate the effectiveness of the data-based Q-CurL in considering the difficulty of the data in training, we consider the scenario of fitting corrupted labels. Given a probability $p$ ($0\leq p\leq 1$) representing the noise level, the label $y_{i}\in\{0,1\}$ of quantum state |$\ket{\psi_{i}}$ is transformed to the label $1-y_{i}$ with probability $p$, while it remains the true label with probability $1-p$. Figure 3 illustrates the performance of trained QCNN on test data across different noise levels. There is no significant difference at low noise levels, but as the noise level increases, the conventional training fails to generalize effectively. In this case, introducing data-based Q-CurL in training (red lines) reduces the test loss and enhances testing accuracy compared to the conventional method (blue lines).
Discussion.— The proposed Q-CurL framework can enhance training convergence and generalization in QML with quantum data. Future research should investigate whether Q-CurL can be designed to improve trainability in QML, particularly by avoiding the barren plateau problem. For instance, curriculum design is not limited to tasks and data but can also involve the progressive design of the loss function. Even when the loss function of the target task, designed for infeasibility in classical simulation to achieve quantum advantage [26], is prone to the barren plateau problem, a well-designed sequence of classically simulable loss functions can be beneficial. Optimizing these functions in a well-structured curriculum before optimizing the main function may significantly improve the trainability and performance of the target task.
@@ Acknowledgements.
The authors acknowledge Koki Chinzei and Yuichi Kamata for fruitful discussions.

### v2

@@ Abstract
Quantum machine learning (QML) requires significant quantum resources to achieve quantum advantage. Research should prioritize both the efficient design of quantum architectures and the development of learning strategies to optimize resource usage. We propose a framework called quantum curriculum learning (Q-CurL) for quantum data, where the curriculum introduces simpler tasks or data to the learning model before progressing to more challenging ones. We define the curriculum criteria based on the data density ratio between tasks to determine the curriculum order. We also implement a dynamic learning schedule to emphasize the significance of quantum data in optimizing the loss function. Empirical evidence shows that Q-CurL significantly enhances the training convergence and the generalization for unitary learning tasks and improves the robustness of quantum phase recognition tasks. Our framework provides a general learning strategy, bringing QML closer to realizing practical advantages.
@@
Introduction.— In the emerging field of quantum computing (QC), there is promise in using quantum computers at sufficiently large scales to solve certain machine learning (ML) problems much more efficiently than classical methods. This synergy between ML and QC has led to the creation of a new field known as quantum machine learning (QML) [1, 2], even though the practicality of its applications remains unclear. There is a question as to whether speed is the only metric by which QML algorithms should be judged [3]. In fact, classical ML focuses on identifying and reproducing features from data statistics. The initial hope is that QML can detect correlations in the classical data or generate new patterns that would be difficult for classical algorithms to achieve [4, 5, 6, 7, 8]. However, there is no clear evidence that analyzing classical data inherently requires quantum effects. This suggests a fundamental shift in the research community’s perspective: it is preferable to use QML on data that is already quantum in nature [9, 10, 11, 12, 13, 14].
@@
The typical learning process in QML involves extensive exploration within the domain landscape of a loss function. This function measures the discrepancy between the quantum model’s predictions and the actual values, aiming to locate its minimum using classical computers. However, the optimization often encounters pitfalls such as getting trapped in local minima [15, 16] or barren plateau regions [17]. These scenarios require substantial quantum computational resources to navigate the loss landscape successfully. Additionally, improving task accuracies necessitates evaluating numerous model configurations, especially against extensive datasets. Given the limitation of current quantum resources, this poses a considerable challenge. Therefore, in designing QML models, we must focus not only on their architectural aspects but also on efficient learning strategies.
@@
Task-based Q-CurL.— We formulate a framework for task-based Q-CurL. The target of learning is to find a function (or hypothesis, prediction model) $h:{\mathcal{X}}\to{\mathcal{Y}}$ within a hypothesis set ${\mathcal{H}}$ that approximates the true function $f$ mapping ${\bm{x}}\in{\mathcal{X}}$ to ${\bm{y}}=f({\bm{x}})\in{\mathcal{Y}}$ such that $h({\bm{x}})\approx f({\bm{x}})$. To evaluate the correctness of the hypothesis $h$ given the data $({\bm{x}},{\bm{y}})$, the loss function $\ell:{\mathcal{Y}}\times{\mathcal{Y}}\to\mathbb{R}$ is used to measure the approximation error $\ell(h({\bm{x}}),{\bm{y}})$ between the prediction $h({\bm{x}})$ and the target ${\bm{y}}$. We aim to find a hypothesis $h\in{\mathcal{H}}$ that minimizes the expected risk over the distribution $P({\mathcal{X}},{\mathcal{Y}})$:
@@
Depending on the problem, we can decide the curriculum weight $c_{M,m}$, where a larger $c_{M,m}$ indicates a greater benefit of solving ${\mathcal{T}}_{m}$ for improving the performance on ${\mathcal{T}}_{M}$. We evaluate the contribution of solving task ${\mathcal{T}}_{i}$ to the main task ${\mathcal{T}}_{M}$ by transforming the expected risk of training ${\mathcal{T}}_{M}$ as follows:
@@
The curriculum weight $c_{M,m}$ can be determined using the density ratio $r({\bm{x}},{\bm{y}})=\dfrac{p^{(M)}({\bm{x}},{\bm{y}})}{p^{(m)}({\bm{x}},{\bm{y}})}$ without requiring the density estimation of $p^{(M)}({\bm{x}},{\bm{y}})$ and $p^{(m)}({\bm{x}},{\bm{y}})$. The key idea is to model $r({\bm{x}},{\bm{y}})$ using a linear model $\hat{r}({\bm{x}},{\bm{y}}):={\bm{\alpha}}^{\top}{\bm{\phi}}({\bm{x}},{\bm{y}})=\sum_{i=1}^{N_{M}}\alpha_{i}\phi_{i}({\bm{x}},{\bm{y}}),$ where the vector of basis functions is ${\bm{\phi}}({\bm{x}},{\bm{y}})=(\phi_{1}({\bm{x}},{\bm{y}}),\ldots,\phi_{N_{M}}({\bm{x}},{\bm{y}}))$, and the parameter vector ${\bm{\alpha}}=(\alpha_{1},\ldots,\alpha_{N_{M}})^{\top}$ is learned from data [23].
@@
Data-based Q-CurL.— We present a form of data-based Q-CurL that dynamically predicts the easiness of each sample at each training epoch, such that easy samples are emphasized with large weights during the early stages of training and conversely. Apart from improving generalization, the benefit of data-based Q-CurL lies in its resistance to noise, which is especially needed in QML. Existing QML models can accurately fit partially corrupted labels to quantum states in the training data but fail on the test data [25]. We show that data-based Q-CurL can enhance the robustness based on the dynamic weighting of the difficulty fitting quantum data to corrupted labels.
@@
To evaluate the effectiveness of the data-based Q-CurL in considering the difficulty of the data in training, we consider the scenario of fitting corrupted labels. Given a probability $p$ ($0\leq p\leq 1$) representing the noise level, the true label $y_{i}\in\{0,1\}$ of training state |$\ket{\psi_{i}}$ is transformed to the label $1-y_{i}$ with probability $p$, while it remains the true label with probability $1-p$. Figure 3 illustrates the performance of trained QCNN on test data across different noise levels. There is no significant difference at low noise levels, but as the noise level increases, the conventional training fails to generalize effectively. In this case, introducing data-based Q-CurL in training (red lines) reduces the test loss and enhances testing accuracy compared to the conventional method (blue lines).
Discussion.— The proposed Q-CurL framework can enhance training convergence and generalization in QML with quantum data. Future research should investigate whether Q-CurL can be designed to improve trainability in QML, particularly by avoiding the barren plateau problem. For instance, curriculum design is not limited to tasks and data but can also involve the progressive design of the loss function. Even when the loss function of the target task, designed for infeasibility in classical simulation to achieve quantum advantage [26, 27], is prone to the barren plateau problem, a well-designed sequence of classically simulable loss functions can be beneficial. Optimizing these functions in a well-structured curriculum before optimizing the main function may significantly improve the trainability and performance of the target task.
@@ Acknowledgements.
The authors acknowledge Koki Chinzei and Yuichi Kamata for their fruitful discussions. Special thanks are extended to Koki Chinzei for his valuable comments on the variations of the Q-CurL game, as detailed in the Supplementary Materials.